Grade 12th passMechanics

From a bmotorboat moving downstream with velocity 2m/s with respect to river, a stone is thrown. The stone falls on an ordinary boat at the instant when the motorboat collides with the ordinary boat. The velocity of the ordinary boat with respect to the river is equal to zero. The river flow velocity is given to be 1m/s. The initial velocity vector of the stone with respect to earth isTake the value of st =10m/s^2. Initial separation between the boats is 20m.

Vaibhav
8 Years agoGrade 12th pass

2 Answers

Shubham
7 Years ago
3i++50j  is the right answer Because the realtive velocity is given 2m/s and the velocity of the  river is 1m/s  .time t=20/2=10sec.
Range=2uu sin¥cos¥/g=100/2=50m/s 
T=2usin¥/gn
 
 
Khimraj
7 Years ago
velocity of boat with respect to the water =Vbw=2m/s   ,flow velocity=1m/s
so velocity of boat A=3m/s
and velocity of boat B=1m/s
velocity of boat A with respect to boat B =2m/s and their relative distance =20
so they will collide in time t
using equation of motion
S=ut+1/2at^2
20=2t+0
t=10
at the same time the stone will collide the boat B
S=ut+1/2 at^2
20=u*10+1/2*10*100
u=-48m/s